# A Unified Causal Framework for Nonlinear Electrodynamics Black Hole from Courant-Hilbert Approach: Thermodynamics and Singularity

### v1

@@ 5.3 Naked singularity of $AdS$-Log black hole
$\mathcal{K}=12\biggl(m-\frac{8\,\,2^{3/4}e^{-3/4\gamma}Q^{3/2}\Gamma(\tfrac{1}{4})\Gamma(\tfrac{5}{4})}{9\pi^{1/2}\lambda^{1/4}}\biggr)^{2}\,r^{-6}+\mathcal{O}(r^{-5}).$ (98)
@@
$r\,f(r)=-m+\frac{8\times 2^{\tfrac{3}{4}}e^{-\tfrac{3}{4}\gamma}Q^{3/2}\Gamma(\tfrac{1}{4})\Gamma(\tfrac{5}{4})}{9\,\pi^{1/2}\lambda^{1/4}}+\Big(k-\frac{2\sqrt{2}e^{-\tfrac{\gamma}{2}}Q}{\sqrt{\lambda}}\Big)\,r+\cdots.$ (99)
@@
$\lambda>4e^{-\gamma}Q^{2}$ (101)
For $\lambda>4e^{-\gamma}Q^{2}$, the spacetime admits at most one horizon. The existence of this horizon is determined by examining the function $rf(r)$, which vanishes at the horizon radius. Furthermore, taking $\lim_{r\to 0}\,r\,f(r)$, we find:
$\lim_{r\to 0}\,r\,f(r)=-m+\frac{8\times 2^{\tfrac{3}{4}}e^{-\tfrac{3}{4}\gamma}Q^{3/2}\Gamma(\tfrac{1}{4})\Gamma(\tfrac{5}{4})}{9\,\pi^{1/2}\lambda^{1/4}}.$ (102)
@@
$m\,<\,\frac{8\times 2^{\tfrac{3}{4}}e^{-\tfrac{3}{4}\gamma}Q^{3/2}\Gamma(\tfrac{1}{4})\Gamma(\tfrac{5}{4})}{9\,\pi^{1/2}\lambda^{1/4}}.$ (103)
@@ 6 Discussion and Conclusions
The work by Russo and Townsend Russo:2025NEDBH provides a novel formulation of self-dual nonlinear electrodynamics where interactions are determined by an auxiliary-field potential. A central achievement is the resolution of a long-standing problem: the derivation of an explicit Lagrangian for the generic ‘analytic’ theory, which is accomplished by restricting to even potential functions. This framework proves powerful for analyzing black hole solutions, yielding new perspectives on their singularities and overall spacetime structure.

### v2

@@ 5.3 Naked singularity of $AdS$-Log black hole
$\mathcal{K}=12\biggl(m-\frac{2\times 2^{\tfrac{3}{4}}e^{-\tfrac{3}{4}\gamma}Q^{3/2}\Gamma(\tfrac{1}{4})^{2}}{9\pi^{1/2}\lambda^{1/4}}\biggr)^{2}\,r^{-6}+\mathcal{O}(r^{-5}).$ (104)
@@
$r\,f(r)=-m+\frac{2\times 2^{\tfrac{3}{4}}e^{-\tfrac{3}{4}\gamma}Q^{3/2}\Gamma(\tfrac{1}{4})^{2}}{9\pi^{1/2}\lambda^{1/4}}+\Big(k-\frac{2\sqrt{2}e^{-\tfrac{\gamma}{2}}Q}{\sqrt{\lambda}}\Big)\,r+\cdots.$ (105)
@@
$\lambda>8e^{-\gamma}Q^{2}$ (107)
For $\lambda>8e^{-\gamma}Q^{2}$, the spacetime admits at most one horizon. The existence of this horizon is determined by examining the function $rf(r)$, which vanishes at the horizon radius. Furthermore, taking $\lim_{r\to 0}\,r\,f(r)$, we find:
$\lim_{r\to 0}\,r\,f(r)=-m+\frac{2\times 2^{\tfrac{3}{4}}e^{-\tfrac{3}{4}\gamma}Q^{3/2}\Gamma(\tfrac{1}{4})^{2}}{9\pi^{1/2}\lambda^{1/4}}.$ (108)
@@
$m\,<\,\frac{2\times 2^{\tfrac{3}{4}}e^{-\tfrac{3}{4}\gamma}Q^{3/2}\Gamma(\tfrac{1}{4})^{2}}{9\pi^{1/2}\lambda^{1/4}}.$ (109)
For the logarithmic theory, we find from the self-energy in Eq. (74) that the condition $m<U^{(0)}$ leads to the formation of a naked singularity.
The panel in Fig. (8(a)) illustrates the phase space of solutions as a function of the charge-to-mass ratio and the mass, with the parameter fixed according to Eqs. (101) and (103) for (GBI) theory. Similar behavior is shown for the logarithmic theory, with its extremal mass plotted via a numerical method in Fig. (8(b)).
Figure 8: Parameter space $(\tilde{Q},m)$ for static Born-Infeld black holes in Fig. 8(a) and for the logarithmic electrodynamics theory in Fig. 8(b). The $\tilde{Q}=e^{-\gamma/2}Q$ is a effective charge. The thick black curve corresponds to Schwarzschild-like black holes, while the region between this curve and the magenta dot-dashed curve represents Reissner-Nordström-like black holes.The magenta dot-dashed curve describes extremal black holes; it terminates at the intersection with the thick black curve-the ‘star point’. Here we have set $k=1$, $\Lambda=0$, and fixed the coupling $\lambda=1$.
@@ 6 Discussion and Conclusions
The work by Russo and Townsend Russo:2025fuc provides a novel formulation of self-dual nonlinear electrodynamics where interactions are determined by an auxiliary-field potential. A central achievement is the resolution of a long-standing problem: the derivation of an explicit Lagrangian for the generic ‘analytic’ theory, which is accomplished by restricting to even potential functions. The Russo and Townsend framework proves powerful for analyzing black hole solutions, yielding new perspectives on their singularities and overall spacetime structure Russo:2025NEDBH.